OpenStudy (anonymous):

Calculate the derivative of the given function without using either the product or quotient rule s(x) = 8 / x^2

OpenStudy (campbell_st):

rewrite the function \[s(x) = 8x^{-2}\] nowapply the normal rule for differentiating

OpenStudy (anonymous):

use the power rule. the function really is 8x^-2.

OpenStudy (anonymous):

-64x?

OpenStudy (campbell_st):

nope... the rule is \[f(x) = x^n..... then.... f'(x) = nx^{n - 1}\]

OpenStudy (campbell_st):

so its \[s'(x) = 8 \times -2 x^{-2 -1}\] just finish it off.

OpenStudy (anonymous):

8 ( -1/8x) is this correct so far?

OpenStudy (campbell_st):

nope... just look at my last posts..

OpenStudy (anonymous):

i did, i dont know how to calculate when i reach \[8(-2x)^-3\]

OpenStudy (campbell_st):

almost you can multiply 8 and -2 to get the coefficient of x then rewrite the problem in index form.

OpenStudy (campbell_st):

opps in non index form

OpenStudy (campbell_st):

\[s'(x) = \frac{-16}{x^3}\]

OpenStudy (campbell_st):

the power of -3 is only operating on x.... and not -2

OpenStudy (anonymous):

so -16x^-3. but im stuck on how to calculate it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i got the concept now, thanks for the help

OpenStudy (campbell_st):

what do need to calculate, you have the derivative. Do you have a point or x-value to substitute into the derivative?